Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
vanhove, mours, romboli, paoli, fabiozzi - 14:19 Thursday 20 November 2025 (68197) Print this report
NEB NCal systematics and survey

20/11/2025

This morning, we performed a geometrical survey of NEB NCals and swapped NEN and NWN rotors. Here below the details.

All times in UTC. Use notation as in wiki page

7:13 : Jigs placed on NE and NW platines between 7:13 and 7:14 (see position plot)

Positions of rotors before survey and changes:

-NWN : rotorbox 2021-06 (wiki table says this box contains R4-17)

-NEN : rotorbox 2022-01 (wiki table says this box contains R4-16)

~8:53 Start calibrating station 1 for measurement of East setup

~9:19:56 : NEF measure; --> fail?

remeasure calibrating points from 9:26:52 to 9:34:28

9:34:28 : NEF point

9:35:05 : NEN point ->9:35:49

9:36:34 : NET point --> 9:37: 16

9:37:50: point 210_aux (to check between stations)

9:42:33 : Start dismounting NCal NEN 

9:56:17 Start calibrating station 2 for measurement of West setup

10:00:07 sphere on NWF point

10:00:38 sphere on  NWN point 

10:01:34-49 move sphere back to NWF

10:02:35: move sphere to NWT --> 10:03:24

measure 210-aux and put rotorbox 2022-01 on NWN

10:08:35 measure NWN_swap; 10:09:22 measure NWF_swap - 10:10:05 measure NWT_swap -->

10:16:29 rotorbox installed on rotorbox 2021-06 on WEN

~10:23:01 start calibrate station 3 for East "swap" measurement

10:25:38 NEF -> 26:13-26:21 NEN -> 26:46-27:14 NET ->27:45-28:04 210_aus 28:54

Remove jigs from East setup (10:29:39-10:30-15)

WEB_pos_sensors.png shows the position measurement of East and West setup during all the today work

11:00 unexpected LV problem for NNN observed: 32V gives 0 current.

  • change it from channel 1 ncal east Power Ssupply to channel 2 ncal south Power Supply
  • It did not fix the problem unfortunately

11:05 we restart all rotors (except NNN which is not spinning) at twice the nominal frequency and leave 

 

 

Images attached to this report
Comments to this report:
a.paoli - 19:43 Thursday 20 November 2025 (68208) Print this report

The result of the NE survey is reported in the entry 68206

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×